### v1

# On AI-Inspired UI-Design

Abstract
Graphical User Interface (or simply UI) is a primary mean of interaction between users and their device. In this paper, we discuss three major complementary approaches on how to use Artificial Intelligence (AI) to support app designers create better, more diverse, and creative UI of mobile apps. First, designers can prompt a Large Language Model (LLM) like GPT to directly generate and adjust one or multiple UIs. Second, a Vision-Language Model (VLM) enables designers to effectively search a large screenshot dataset, e.g. from apps published in app stores. The third approach is to train a Diffusion Model (DM) specifically designed to generate app UIs as inspirational images. We discuss how AI should be used, in general, to inspire and assist creative app design rather than automating it.
The UI of apps plays a pivotal role in the overall user experience, as it translates complex software functions into visually intuitive elements that users can easily understand and navigate. Chen et al. demonstrated in a recent study that well-designed UI can affect user satisfaction and distinguish an app from its competitors [Chen:HowShouldImprove:2021]. A good UI should be functional, engaging, and aesthetically pleasing. Achieving this level of design requires a high degree of creativity [Pham:SEEM:18].
The typical design process of the UI of an app involves several key activities: setting design goals, conducting user research, identifying main features and usage scenarios, wireframing, visual design, prototyping, development hand-off, and gathering user feedback. Commercial tools such as Figma [^1] https://www.figma.com/ and Sketch [^2] https://www.sketch.com/ provide basic templates and UI components, which serve as valuable starting points for wireframing, visual design, and prototyping. However, designers often require additional inspiration to effectively design complex app features and corresponding UIs. To gather inspiration, app designers frequently explore existing UIs, e.g. on their mobile phones, on app stores such as Google Play and Apple App Store, or on design portfolio platforms like Mobbin [^3] https://mobbin.com/ and Dribbble [^4] https://dribbble.com/.
Recent advancement in AI, particularly Generative AI and the availability of large foundation models is revolutionising the process of app development in general and app UI design in particular. Recent studies reveal that AI can be effectively used for accurate text-to-UI retrieval [Wei:GUingMobileGUI:2024, Kolthoff:DatadrivenPrototypingNaturallanguagebased:2023]. Moreover, Generative AI, encompassing both image and text generation, can be applied to UI generation, offering an infinite source of design inspiration [Wei:BoostingGUIPrototyping:2023, Feng:DesigningLanguageWireframing:2023]. Several commercial tools are already available for text-to-UI generation, such as Uizard’s Autodesigner [^5] https://uizard.io/autodesigner/, Galileo AI [^6] https://www.usegalileo.ai/explore, Visily’s AI [^7] https://www.visily.ai/ai-ui-design-generator/, and JS.Design’s AI [^8] https://jsai.cc/ai/create. Many app development teams are starting to explore these trends. The central question is thereby: how can AI assist a good UI design, or perhaps even automate it?
Figure 1: Overview of an AI-inspired Design process for UI-intensive software (particularly apps) inspired by [Gohar:TR:2024, HBR:2024].
Boosting Creativity with AI
In a recent fascinating study on creativity in general, Gohar and Utley investigated the impact of using ChatGPT on ideation and creative problem solving [Gohar:TR:2024]. The authors asked practitioner teams from different companies to engage in creative solving tasks for problems from their organisations. They then asked product owners to assess the resulting solutions ideas and asked the teams for assessing their experiences. The authors found that the teams who used ChatGPT created more ideas compared to those who did not. However the effect was rather small (only 8% increase). The authors also observed that Generative AI was “helping people develop fewer truly bad ideas but AI assistance was also leading to more average ideas”. They concluded that if teams want to outperform in creative problem-solving using AI, they need to adhere to certain practices but most teams are not [HBR:2024].
What is fascinating about this study is that it shows that AI can give the impression to teams that they can “perform fast” when they simply delegate the work to AI tools. In reality, this can be counterproductive as it stimulate human laziness. The authors suggest a simple process which they call FIXIT to guide a AI-supported problem solving, particularly highlighting that AI should be used in a conversational iterative why to get the best of human creativity, instead of transactional “do the work for me” manner.
In Figure 1 we adopt and extend Gohar and Utley’s findings into the context of app design. We suggest an AI-inspired app design, including six steps:
  1. 1.
    First, app designers should scope the app requirements as much as possible. Starting from an overall aim or purpose of the app, a list of features, sub-features, and possible user stories can, e.g., be created by prompting an LLM [Nakagawa:MAPEKLoopbasedGoal:2023].
  2. 2.
    Once a preliminary “backlog” of features or user stories is available (in text form), developers and designers can engage in several ideation steps, individually, supported by an AI tool, and in team. Research has repeatedly shown the importance of balancing between individual ideation and join brainstorming sessions for creative work. Therefore, it is crucial that designers take time by themselves in the second step to think about the target app feature, user interface, and user experience. This can be brief, according to Gohar and Utley about 15-30 minutes.
  3. 3.
    Third, to get the most fitting and inspiring results from foundation models, a prompt or a query should be as detailed and with as much context as possible. This context can, e.g. be provided to an LLM in form of a problem statement, notes, the elicited features from step 1, or even some UI sketches and previous screens from the app.
  4. 4.
    The fourth step is the core of AI-inspired design. A human designer should engage in a conversation with AI including multiple queries or prompts, assuming that the best results are not necessarily retrieved straight away. For instance, when prompting ChatGPT for ideas, a designer can immediately ask for additional recommendations, ask other LLMs to review the recommendations or review by herself and explain to ChatGPT why certain recommendations are good or bad. This step requires careful critical thinking about the AI recommended results. Critical questions can be what UIs or design elements make most sense and why? What are most interesting and intuitive designs and why?
  5. 5.
    The fifth step focuses on the group or team setting, ideally seeking for other perspectives such as a customer, a lead user, a domain or a technology expert. This step can follow typical brainstorming or design thinking processes [Pham:SEEM:18]. The team can also use AI in parallel to seek for feedback or get additional suggestions. At the end, the team should decide about the design to follow.
  6. 6.
    Finally, the design and the corresponding logic will be implemented, tested, and released to users. In some cases, AI-generated UI code can also be reused.
In the following we present three state-of-the-art approaches on how to get UI design recommendations from foundation models: 1) by generating app UIs by prompting Large Language Models (LLM); 2) by leveraging Vision-Language Models (VLM) to search large screenshot repositories; 3) by training a diffusion model (DM) to generate creative app screens. Readers can try out these approaches in our accompanying online material [^9] https://github.com/Jl-wei/ai-gen-ui and https://github.com/Jl-wei/guing. We use a health monitoring app as illustrative example as shown on Table Boosting Creativity with AI.
Table 1: Examples of UIs obtained using three different approaches with the input “Health Monitoring Report”
{tblr}
[…]
UI Generation with LLMs [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UI Retrieval with VLMs [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UI Generation with DMs [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[1pt]
UI Generation with LLMs
Large Language Models (LLMs), such as GPT, represent a category of AI models designed for interpreting, generating, and manipulating human language. These models, characterised by their ability to generate human-like text based on extensive training data, have shown remarkable proficiency in a wide range of applications, including automated content creation, customer service, and code generation. LLMs are based on deep learning techniques, particularly neural networks with many layers, which accounts for the "large" in their name. LLMs are pre-trained through “next-word prediction” on vast amounts of text data from diverse sources, such as books, articles, websites, and other text repositories, which allows them to understand a wide range of topics [Zhao:SurveyLargeLanguage:2023].
Given an app page description, it is possible to automatically generate UIs, e.g. HTML code, by prompting large language models. In a recent study, Fend et al. fine-tuned GPT-3.5 model for UI generation [Feng:DesigningLanguageWireframing:2023]. As expected, our preliminary assessment indicates that GPT-4 can achieve excellent results even without fine-tuning. We thus focus on discussing GPT-4 in the following.
Figure 2: Overview of UI Generation with LLMs
The process of UI generation using LLM, illustrated in Figure 2, includes three key steps:
Description Refinement
In this step, the LLM refines the high-level feature or page description into a list of detailed UI sections. For example, given “health monitoring report” as input, it will refine it to a list of UI sections, such as “header section”, “profile section”, “summary section”, “charts section”. Each section will include a description and a corresponding code example. This refinement ensures that the generated HTML is more detailed and structured.
HTML Generation
In this step, the LLM uses detailed UI sections as input to generate the corresponding HTML code. Given the complexity of code generation, this task necessitates an advanced LLM to increase the likelihood of producing high-quality HTML output. Our preliminary assessment indicates that GPT-4 significantly outperforms GPT-3.5 in this regard.
HTML Adjustment
The generated HTML code might lack essential UI elements, such as a footer, or have incorrect image sizes. This step involves adapting the code with additional prompts to address these issues. For example, by asking GPT-4 to “add the footer” to the provided HTML code, it will include the “footer section” in the HTML, as illustrated in Figure 2.
Pros
  • Reusable code: The output of this approach is HTML code, which can be partially or fully reused in subsequent development. The generated code can be easily modified using additional prompts.
  • Low hardware requirements with cloud LLMs: The hardware requirements for deploying this approach are minimal when using a cloud-deployed LLM such as GPT-4.
Cons
  • Aesthetic issues: Occasionally, the generated HTML exhibits alignment issues, negatively impacting its aesthetics. Sometimes, the generated HTML might lack CSS styling altogether.
  • No images: Although HTML code generated by the LLMs include icons provided by Material 3 library [^10] https://m3.material.io/styles/icons/overview, LLMs are incapable of generating images. This limitation may constrain the diversity of UIs created using this method.
  • High latency: In our preliminary assessment, it takes more than 30s to generate a UI with GPT-4o. This can be mitigated by parallelising multi generations. However, each individual generation still requires a minimum of 30 seconds.
  • Privacy issues with cloud LLMs Utilizing cloud LLMs like GPT-4 or Claude-3111111https://www.anthropic.com/news/claude-3-family raises data privacy concerns, as it necessitates uploading your data to the cloud.
  • High hardware requirement with local LLMs: Deploying a local LLM, such as Llama 3121212https://llama.meta.com/llama3/, requires expensive hardware, making it a less viable option for many users.
Keep in mind
  • Tuning the temperature of LLMs: The temperature of the LLM has a clear impact on the diversity of the generated UI. To get a variety of designs and get inspiration, avoid setting the temperature to 0.
UI Retrieval with VLMs
Instead of generating new UIs, an alternative approach is to search for existing UIs in other apps using textual queries. The most straightforward method for retrieving UIs is searching app stores like Google Play or Apple App Store. However, this approach operates at the app level (i.e. the text describing the app) rather than the UI level. To tackle this problem, Kolthoff et al. proposed RaWi [Kolthoff:DatadrivenPrototypingNaturallanguagebased:2023] an approach using text embedding (from metadata available about the UI) for accurate UI retrieval. Our recent work demonstrates that VLMs clearly surpasses text embedding model in text-to-UI retrieval [Wei:GUingMobileGUI:2024].
Figure 3: Overview of UI Retrieval with VLMs (adapted from [Wei:GUingMobileGUI:2024])
Vision-Language Models, such as CLIP [Radford:LearningTransferableVisual:2021], are multimodal models, capable of learning from both images and text (i.e. multiple modalities). These models are pre-trained on large-scale image-caption datasets using contrastive learning techniques. VLMs have the capability to convert images and text into a shared multimodal embedding space, ensuring that semantically similar images and texts are closely aligned. By computing the similarity between images and textual queries, these models can be effectively used in various tasks, such as zero-shot image classification (e.g., iOS vs. Android Design) and text-to-image retrieval. Text-to-image UI retrieval with VLMs requires four steps.
VLM Fine-tuning
The vanilla CLIP model [Radford:LearningTransferableVisual:2021], originally trained on general image-caption datasets that include images of dogs, buildings, and other common subjects, may not perform well in text-to-UI retrieval task. To achieve better performance, fine-tuning the VLM with a UI-caption dataset, such as GPSCap [Wei:GUingMobileGUI:2024], is recommended. Instead of fine-tuning your own VLM model for tasks, we have made our UIClip model publicly available [^13] https://huggingface.co/Jl-wei/uiclip-vit-base-patch32.
UI Repository Preparation
As the UI retrieval is to perform searching on an existing UI dataset, the first step is to prepare a UI repository. There are existing large UI datasets, like Rico [Deka:RicoMobileApp:2017] and GPSRepo [Wei:GUingMobileGUI:2024]. Besides, users can create their own repository, e.g. though automatic UI exploration [Chen:AutomaticallyDistillingStoryboard:2023].
UI Image Processing
UI retrieval using VLM is achieved by calculating the similarity between the query embedding and the embeddings of the UIs. Thus, pre-computing the embeddings of the images in the UI repository is essential. As shown in Figure 3, all images in the UI repository are converted into embeddings using the image encoder of the VLM. These embeddings are then stored in a vector database like Faiss [^14] https://github.com/facebookresearch/faiss for subsequent searches.
UI Image Querying
While using the UI search engine, the user query is converted into a text embedding using the text encoder of the VLM. The resulting text embedding is compared with image embeddings stored in the vector database, and the top-k most similar images are selected as output.
Pros
  • High quality UI images: The UI images originates from existing apps. Thus, their quality is usually much better than those generated by an AI model.
  • Low latency: The UI retrieval approach takes less than 1 second to obtain results. It is significantly faster than UI generation.
  • Link to source app: Each retrieved UI image is linked to its source app. This enables practitioners to explore the source app for implementation details and e.g. gather user feedback on its design.
Cons
  • Limited diversity: Unlike the UI generation which leverages randomness to produce unlimited number of UIs, the UI retrieval approach recommend UIs from existing apps, thereby limiting diversity. It is therefore important to explore loosely related apps and designs by variating the queries.
  • Irrelevant results: Some retrieved UIs may be entirely irrelevant to the given query.
  • Complexity of deployment: Deploying an own UI search engine necessitates several hundred gigabytes of disk space to store a large UI repository. Additionally, maintaining a vector database is necessary for its operation.
Keep in mind
  • Copyright The retrieved UIs cannot be reused directly in products due to likely copyright restrictions.
UI Generation with DMs
Diffusion models (DMs), such as Stable Diffusion [Rombach:HighResolutionImageSynthesis:2022], are widely used in text-to-image generation. These models operate by iteratively refining an initial random noise image through a series of denoising steps guided by the input text. The core principle behind diffusion models is to reverse the diffusion process, whereby an image is progressively degraded into noise. By training on a large images-caption dataset, the model learns to generate coherent and visually appealing images that match the given textual input.
Using text-to-image models, it is possible to generate UI images from app page descriptions. Wei et al. recently introduced UI-Diffuser [Wei:BoostingGUIPrototyping:2023], a UI image generator built on Stable Diffusion [Rombach:HighResolutionImageSynthesis:2022]. However, UI-Diffuser necessitates both a page description and the image layout information to achieve reasonable performance. With this paper, we release UI-Diffuser-V2151515https://huggingface.co/Jl-wei/ui-diffuser-v2, which generates relevant UI images using only the page description of the apps.
The UI generation can be achieved via cloud-based image generator, such as Midjourney [^16] https://www.midjourney.com/home and DALL·E 3171717https://openai.com/index/dall-e-3/. However, these generators are not specifically trained for creating mobile UIs, often resulting in skewed UI designs, and suffer from higher latency. For optimal performance, fine-tuning a diffusion model with a UI-caption dataset is recommended.
We fine-tuned the Stable Diffusion model with the recently available GPSCap dataset [Wei:GUingMobileGUI:2024]. GPSCap (Google Play Screenshot Caption) is a large dataset with 135k UI-caption pairs. It was created by automatically performing classification, cropping, and caption extraction from the app introduction images of Google Play. We trained the Stable Diffusion model for five epochs on a machine equipped with an NVIDIA T4 GPU, resulting in the UI-Diffuser-V2 model.
The model is straightforward to use: by inputting a page description, it generates corresponding UI images. It is also capable of producing multiple images in one generation, though this requires additional VRAM. Our preliminary assessments indicate that the NVIDIA T4 can generate a batch of over 20 images simultaneously.
Pros
  • High diversity: This approach can generate highly diverse UI images with varying layouts, providing an excellent source of inspiration.
Cons
  • Aesthetic issues: There will likely be graphical issues with the UI generated with DM, such as unreadable text and misaligned components. In some instances, the DM generate significantly flawed images.
  • Limited reusability: The aforementioned aesthetic problems, coupled with the intrinsic nature of images, greatly hinder the immediate reusability of UIs created with DM.
Keep in mind
  • Tuning the batch size Batch size refers to the number of images obtained in a single generation. Using the NVIDIA T4, it is capable of generating a UI within 3 seconds when the batch size is set to 1. When the batch size is increased to 20, the generation time is approximately 30 seconds.
Conclusion
Foundation models, particularly LLMs, bear a great potential to revolutionise software development by generating various artefacts including source code, documentation or test cases. One particular artefact that usually requires creativity and multiple iteration of tuning is the design of the UI. We discussed how to use large language models, diffusion models, vision-language models to retrieve and generate relevant, diverse and inspiring designs .
Each approach has pros and cons. None of them is perfect. In practice, app teams have to carefully choose or combine multiple approaches. Our assessment is preliminary and based on related work and a running example shown on Table Boosting Creativity with AI. More research is needed to gain evidence on how to combine the approaches as well as important factors to consider such as team size, domain, novelty of the features, or designers skills. Additionally, while these AI techniques can significantly enhance creativity by serving as sources of inspiration, they cannot replace human creativity and experience. Human (i.e. designer and developer) involvement remains essential to the design process [HBR:2024, Andersen:DesignPatternsMachine:2024].
[…]